Your bank pays an interest rate quoted as 4.0% per annum compounded semi-annually. You invest $10,000 into the bank now for a 5 year period. What will your balance be after 5 years?
Answer #1

We use the formula:
A=P(1+r/2)^2n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=10,000*(1+0.04/2)^(2*5)

=10,000*1.21899442

=$12189.94(Approx).